DISSOCIATIVE RECOMBINATION OF D_{3}S^{+}: ABSOLUTE CROSS SECTION, BRANCHING FRACTIONS AND ASTROPHYSICAL APPLICATION

\underline{MAGDALENA KAMINSKA}, Institute of Physics, Świetokrzyska Academy, Świetokrzyska 15, PL-25406, Kielce, Poland; ERIK VIGREN, VITALI ZHAUNERCHYK, WOLF D. GEPPERT, MATH-IAS DANIELSSON, MATHIAS HAMBERG, RICHARD D. THOMAS, MATS LARSSON, Department of Physics, Albanova University Center, Stockholm University, SE-106 91, Stockholm, Sweden; MAGNUS AF UGGLAS, Manne Siegbahn Laboratory, Freskativägen 24, SE-104 05, Stockholm, Sweden; and JACEK SEMANIAK, Institute of Physics, Świetokrzyska Academy, Świetokrzyska 15, PL-25406, Kielce, Poland.

The chemistry of sulfur bearing species in molecular clouds is a very important but still not fully explored field. One of the crucial processes occurring there is a dissociative recombination (DR) of H_{3}S^{+}, which is one of the suggested production pathways for H_{2}S. The DR of D_{3}S^{+} has been studied at the heavy ion storage ring CRYRING in Stockholm, Sweden. It has been observed that the main production channel in this process leads to the formation of DS + 2D, whereas D_{2}S + D is only a minor pathway. The obtained value for the reaction rate coefficient is 2.8×10^{−7} at 300 K. We used the obtained data to investigate their effect on the predictions of the gas phase and gas-grain model calculations of the sulfur chemistry and deuterium fractionations in dark clouds employing the UMIST code. The results obtained agree well with observations for abundances and deuterium fractionations for several interstellar sulfur compounds. However, large discrepancies are found for H_{2}S and HCS^{+}, which are grossly underestimated in both models. We conclude that H_{2}S is formed mainly on grain surfaces and not through gas-phase reactions in dark clouds.